Add or substract as indicated. Express your result in simplest form.
3/x^2+4x+3 - 1/x^2-9
Thank you.

Answer by sdmmadam@yahoo.com(530)

3/x^2+4x+3 - 1/x^2-9
=(3/x^2-1/x^2)+4x+(3-9) (grouping like terms)
=(3t-t)+4x-6 where t =1/x^2
=2t+4x-6
=(2/x^2)+4x-6